The velocity of an object is given by the expression v(t) = 3.00 m/s + (4.00 m/s3)t2, where t is in seconds. Determine the position of the object as a function of time if it is located at x = 1.00 m at time t = 0.00 s .




1
Expert's answer
2021-05-25T10:10:34-0400

velocity of an object is given by the expression

"v(t) = 3.00 \;m/s + (3.00\; m/s^3)t^2"

located at x = 1.00 m

at time t = 0.000

Find the position of the object as a function of time.

Now,

"v(t) = 3.00 \;m/s + (3.00 \;m/s^3)t^2 \\

v=\frac{dx}{dt}=3.00 \;m/s + (3.00 \;m/s^3)t^2 \\

\int dx=\int (3.00 m/s + (3.00\; m/s^3)t^2 )dt \\

x=3t+\frac{3t^3}{3}+C \\

t=0 \\

x=1 \\

1 = 3 \times 0 + \frac{3 \times 0^3}{3} + C \\

C=1 \\

x=3t+\frac{3t^3}{3}+1"

The position of the object as a function of time

"x=t^3+3t+1"